An investment cost N800,000. You know that it will produce an inflow after operating cost of N170,000 a year for 10 years. If the opportunity cost of capital is 14 percent, what is the net percent value of the investment? What will the investment be worth at the end of 5years


1
Expert's answer
2021-11-15T17:51:48-0500

The present value of the 10-year stream of cash inflows is:

PV=$"170000 \times (\frac{1}{0.14}- \frac{1}{0.14\times(1.14)^10})=886739.66"


Thus:

NPV = –$800,000 + $886,739.66 = +$86,739.66


At the end of five years, the factory’s value will be the present value of the five remaining $170,000 cash flows:

PV=$"170000 \times (\frac{1}{0.14}- \frac{1}{0.14\times(1.14)^5})=583623.76"